OpenStudy (anonymous):

15 articles are bought some cost 5 cents each and the remainder cost 8 cents each the total amount spent was 9 cents how many of each are bought

OpenStudy (anonymous):

anyone can help me with this question

OpenStudy (anonymous):

Pls check if the total amount spent is 9 cents or ?????

OpenStudy (anonymous):

?????? 15 articles have to cost more than 9 cents....

OpenStudy (anonymous):

its 19

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let the no of 5 cent articles = x then no of 8 cent articles = (15-x) ---- becoz total no of articles is 15 cost of x no of articles of cost 5 cent each = 5*x = 5x cents cost of (15 - x) articles of cost 8 cent each = 8*(15 - x) = (120 - 8x) cents ATQ 5x + (120 - 8x) = 19 5x - 8x = 19 - 120 -3x = -101 x = -101/-3 ?????? Pls recheck your entire question as with given figures it can't be solved..... pls check ALL numbers/figures and revert...

OpenStudy (anonymous):

oh okay thats abit confusin

OpenStudy (anonymous):

total spent should be 90 cents i guess, so that the answer is: 5a+8b=90 a+b=15 a&b are the numbers of articles respectively a=10, b=5

OpenStudy (anonymous):

Harkirat's method is right. He is right that the figures supplied are wrong. Maybe what xwang has suggested might be the right figure.